Why use the k:1 in the problem, find the ratio in which the y- axis divides the line segment joining the points (5,-6) and (-1,-4). Also find the point of intersection?

Grade:10

1 Answers

Arun
5 years ago
Let the coordinates of the point be (0,y) and the y axis divide the line segment in the ratio k:1
(0,y)={(-k+5)/k+1 , (-4k-6)/k+1}
=>0=-k+5/k+1
=>0=-k+5
=>k=5
y=-4k-6/k+1
=>y=-4(5)-6/5+1
=>y=-20-6/6
=>y=-26/6
thus the coordinates of the point is (0,-26/6)